OpenStudy (anonymous):

the surface area of a rectangular prism is found using the formula SA=2(lw+lh+wh) where l is length, w is width, and h is height. if the surface area of a rectangle prism is 324m2, with length 5 m and width 12 m what is the height in meters?

OpenStudy (anonymous):

You first have to plug everything in....324 m^2 is the SA so 324 m^2 = 2[(5m*12m) + (5m*h) + (12m*h)] You want to find h so you can first start by dividing 342 by 2 which will make your equation simpler to solve. So you should have 162 m^2 = (5m*12m) + (5m*h) + (12m*h) Multiply your "lw" together 5m*12m=60m^2 and you have 162 m^2=60 m^2 + (5m*h) + (12m*h) You can subtract 162 m^2 by 60 m^2 leaving 102 m^2 = (5m*h) + (12m*h) Then combine your like terms of h....5m*h + 12m*h = 17m*h so you now have 102m^2 = 17m*h and to get h by itself to solve for it you can then divide by 17 leaving 6m=h making your height equal 6 meters. I left the meters in the equation the whole time so you could see how they cancel each other out making m^2 go back to just m. I took this step by step so you could understand it all, sorry it is long but I hope this helps you.

OpenStudy (anonymous):

Thank you!!!:)

OpenStudy (anonymous):

No problem :]

OpenStudy (anonymous):

@knstagge can you help me with my problem on my profile